# Hyman et al. 2025

### v1

@@ 1 Introduction
In this study, we apply a regression framework to forecast harvest for the recreational sector component of the Gulf gag (Mycteroperca microlepis) stock. Gulf gag are highly targeted by recreational anglers (Hyman2024Effort, ) and the gag fishery is currently in a rebuilding plan (Amendment56, ). Recreational catch-per-unit-effort (CPUE) has declined in recent years (GAGSEDAR72, ) and stock status is currently considered impaired (i.e., overfishing is occurring and the stock is overfished; GAGSEDAR72, ; Gag_INTERIM, ; Amendment56, ). In response, the Gulf of Mexico Fishery Management Council (GMFMC) reduced the recreational gag annual catch limit (ACL) by 80% and reduced the season duration from 214 days in 2022 to 49 days in 2023 (Amendment56, ) and 15 days in 2024 (NOAA2024GagSeason, ). Projections of season duration based on harvest rates from previous years underestimated harvest in the substantially shorter season, and the ACL was exceeded in both 2023 (Amendment56, ; NOAA2024GagSeason, ) and with preliminary 2024 estimates, suggesting intensified fishing effort within the shortened time frame. Consequently, more precise projection methodologies that (1) incorporate changes in harvest rates driven by temporal restrictions (i.e., effort compression) and environmental conditions and (2) quantify the risk of exceeding the harvest quota could enhance future fisheries management decisions. We developed a set of statistical models to generate monthly estimates of harvest rates (pounds per day open) using management, social, environmental, and seasonal predictors. We also demonstrate this framework may be used to generate pre-season forecasts and in-season updates of when a hypothetical quota will be met.
@@ 2.1 Data sources
We compiled recreational gag harvest data for the for-hire sector from the National Marine Fisheries Service (NMFS) Marine Recreational Information Program (MRIP) For-Hire Telephone Survey (FHTS) combined with data from the Angler Point Access Intercept Survey (APAIS) and private recreational data from the Florida Fish and Wildlife (FWC) State Reef Fish Survey (SRFS). This approach aligns with the data sources currently used for gag management, as both the Council and NOAA NMFS monitor recreational gag landings using SRFS and FHTS/APAIS data.
@@
The State Reef Fish Survey (SRFS) was established in 2015 (originally named the Gulf Reef Fish Survey [GRFS]) to improve harvest estimates for offshore species that are less frequently targeted and caught. The survey focuses on reef-associated species, including gag, red grouper (Epinephelus morio), gray triggerfish (Balistes capriscus), red snapper, vermilion snapper (Rhomboplites aurorubens) greater amberjack (Seriola dumerili), hogfish (Lachnolaimus maximus), mutton snapper (Lutjanus analis), and yellowtail snapper (Ocyurus chrysurus). The SRFS estimates CPUE using a combination of data collected on these reef-associated species from the APAIS and data collected from SRFS intercepts, which are conducted at public boat ramps where offshore fishermen are more commonly encountered. The SRFS mail survey is sent to a stratified, random sample of anglers with the free state reef fish angler designation on their Florida saltwater license, asking them to report their fishing activity over the past month. The combination of increased dockside sampling at offshore-access points, a larger sample size, and a more targeted mail survey frame allows the SRFS to produce more precise harvest estimates for reef fish compared to the MRIP survey, which has a broader scope of all saltwater fish species.
@@
Both SRFS and FHTS use dockside sampling to collect catch-per-trip data. In these interviews, trained samplers record details from anglers about their completed fishing trips, including the number and species of fish caught, the size of landed fish, and fishing effort (e.g., hours fished). These catch data are then combined with effort estimates from the mail or telephone surveys to calculate total harvest. Statistical methods are applied to account for sampling variability and potential reporting biases, ensuring robust estimates for fisheries management. To standardize harvest rates from SRFS and FHTS in Florida, we aggregated monthly gag harvest estimates across months with varying open seasons. We then divided each monthly estimate by the number of days open to fishing to calculate daily harvest rates. To avoid division by zero in months when harvest was closed, we added one additional day to the denominator.
@@ 2.2 Predictors considered
We considered multiple regulatory, social, and environmental predictors to model patterns in recreational gag harvest rates (Table 1). Fundamentally, we expected gag harvest rates to be positively related to whether a month was open to gag harvest, as anglers would presumably target gag when the season is open (Hyman2025Catch, ). However, gag temporal regulations along the Gulf Coast of Florida are complex and have historically differed spatially, requiring careful consideration of how to translate these regulations into relevant predictors. From 2012 to 2022, a four-county special season for gag in Florida applied to Franklin, Wakulla, Jefferson, and Taylor counties. During this period, state waters in these counties were opened for recreational gag harvest from April 1 through June 1 or June 30, depending on the year (Hyman2024Effort, ). This special season allowed for local adjustments to manage the species effectively and support angling opportunities. For all Gulf state and federal waters, gag harvest was generally open from June 1 to December 31. To account for spatial differences in gag temporal restrictions, we created three dummy variables denoting whether the four-county season (denoted “$\text{Special}_{Gag}$”) or the remainder of Gulf state and federal waters (denoted “$\text{Full}_{Gag}$”) was open, or neither (denoted “$\text{Closed}_{Gag}$”). Furthermore, for the hurdle component of our regression model (see Analysis section below), we introduced a fourth binary variable, “$\text{Open}_{Gag}$”, which indicated whether either season was open. This was based on the expectation that the probability of observing zero gag harvested per day would depend on whether gag harvest anywhere was permitted during that period.
@@ 2.3 Analysis
where $y_{t}$ denotes the weight of gag harvested per day open in the $t^{th}$ month-year. Meanwhile, $\mu_{t}$ and $\sigma^{2}_{t}$ denote the mean and variance of the Gamma distribution conditioned on $y_{t}>0$, respectively, and $\theta_{t}$ denotes the binomial probability of observing 0 gag caught in month-year $t$. We use $x$, $z$, and $k$ to refer to predictors included in modeling $\mu_{t}$, $\sigma^{2}_{t}$, and $\theta_{t}$, respectively. Similarly, coefficients corresponding to $\mu_{t}$, $\sigma^{2}_{t}$, and $\theta_{t}$ are denoted $\beta$, $\rho$, and $\xi$, respectively. Both $\mu_{t}$ and $\sigma^{2}_{t}$ are related to their predictors and associated coefficients through a log-link function, whereas $\theta_{t}$ is related to its predictors and associated coefficients through a logit-link function. All regression coefficients were assigned relatively uninformative, normally-distributed priors with mean 0 and variance 16.
We used the Stan probabilistic programming language for Bayesian modeling (gelman2015stan, ), accessed via R (Rcore2024, ). Each model involved four parallel Markov chains with 5,000 iterations for warm-up and another 5,000 for posterior sampling, yielding 20,000 total draws for inference. Covariates were deemed significant if their 80% confidence intervals excluded 0 (ipccAR6, ; kruschke2021bayesian, ; van2021bayesian, ).
@@ 2.3.2 Cross-validation
We employed extensive cross-validation (CV) to assess the predictive power of our best-performing statistical model. We considered two CV exercises – block-CV and 1-step-ahead-performance CV (1-SAP CV) – to separately assess the suitability of our model for pre-season and in-season projections, respectively. For our block-CV exercise, we refit the model to truncated datasets eight times – each with one year withheld – and subsequently compared predictions derived from our models to withheld, observed response variables for all months in the withheld year. As managers would only have information from previous years when making pre-season projections, this cross-validation method was a reasonable approach to assessing model pre-season posterior prediction accuracy. In contrast, nowcasting entails making in-season projections for the current time-step based on previous information both from the current year and previous years (carter2015nowcasting, ). As a result, for 1-SAP CV, for each year, we subsequently refit the model 11 times (once for each month). For each refit, we included months 1 to $m$ in our training set and then compared posterior predictive distributions to observed data in month $m+1$.
@@ 2.4 Comparison to current methods
After establishing the posterior predictive accuracy of our models, we conducted simulations to predict the 2025 gag season duration using 1,000 posterior scans from our best-fitting model. To begin, we calculated the adjusted ACT for 2025 by accounting for the 2024 overage penalty—defined as the amount of gag harvested in 2024 (249,000 lbs) that exceeded the 2024 ACL (163,000 lbs). This overage was subtracted from the 2025 ACT (319,000 lbs) specified in the rebuilding plan (Amendment56, ; NOAA2024GagSeason, ; ECFR2025, ):
@@
To develop our simulations, we constructed a design matrix where all predictors, except for gag temporal regulations and seasonal variables, were fixed at their 2024 values. We began by setting the season duration to $d=1$ day, with a starting date of September 1st. Months without any days open to harvest were assigned the “Closed” temporal regulation treatment, whereas months with days open were assigned the “Statewide” treatment. Using our best-fitting model, we simulated posterior-predictive distributions of monthly harvest rates for 2025 across 1,000 posterior scans ($s$). For each scan, monthly harvest rates were multiplied by the number of days open to harvest within the corresponding month to derive posterior predictions of monthly harvest. To estimate total harvest for each scan ($s$), we summed harvest estimates across all months. The process was repeated iteratively, incrementing the season duration ($d$) by one day at a time, until $d=121$, which corresponded to the maximum possible season duration ending December 31st. For each season duration ($d=1,\dots,121$), we calculated the median and 80% posterior predictive interval of total harvest. Additionally, for each season duration, we estimated the probability that the ACT would be met or exceeded as the proportion of posterior predictive scans whose total harvest met or exceeded the ACT. Finally, we compared these projections to those using the historic average harvest rate for September.
@@ 3.1 Model selection and diagnostics
Time-series plots comparing observed and predicted values, along with posterior predictive checks, indicated that model $g_{2}$ provided a good fit to the observed data. The uncertainty in monthly estimates aligned closely with in-sample estimates of monthly gag harvest rates (Fig. 1a). Additionally, diagnostic scatter-plots of median predicted versus observed gag harvest demonstrated satisfactory model performance (Figs. 1b and 1c).
@@ 3.2 Predictors of monthly gag harvest rates
Gag temporal regulations, the annual trend, and seasonal harmonic terms all significantly explained variation in mean (nonzero) recreational gag harvest rates (Table 3.2). Conditioned on a positive outcome and relative to a closed season, months when gag was open to the entire Florida Gulf Coast (FullGag) resulted in higher harvest rates, while months open only for the four-county season (SpecialGag) resulted in lower harvest rates. The duration of the gag season had a clear negative effect on harvest rates, such that the harvest rate was lower during longer gag seasons (Fig. 2) The annual trend was positive and significant, indicating that after accounting for other predictors, gag harvest rates appeared to be increasing over time. Notably, after accounting for the effects of other predictors, past harvest did not significantly explain residual variation in gag harvest rates.
@@
The hurdle parameter, $\theta$ was strongly negatively impacted by whether gag was open to harvest (either in the four-county region or across the Gulf Coast of Florida). When the season was completely closed to gag harvest, the probability of observing 0 gag harvested in a given month was 0.90 (80% CI: 0.84 – 0.95). However, this probability decreased to nearly 0 when gag was open to harvest. Meanwhile, both sinusoidal harmonic terms significantly explained $\sigma^{2}$, suggesting that model uncertainty varied seasonally.
@@ 4.1 Predictors of gag harvest
A key finding of this study is the pronounced effort compression observed in the recreational Gulf gag fishery. While previous research did not identify a relationship between gag season duration and recreational reef fishing effort (e.g., Hyman2024Effort, ), our model revealed a strongly negative effect on gag harvest rates, which increased non-linearly as season duration decreased. As noted by Hyman2024Effort , a lack of contrast in season duration previously may have limited the detection of this effect. The recreational gag fishing seasons have only recently been significantly shortened, with the 2023 and 2024 seasons being the shortest on record. This is supported by our block-CV results: when 2024 data were included in model training, the model accurately predicted effort compression and correctly estimated withheld gag harvest rates for 2023 (Fig. 3). However, when 2024 data were excluded from the training dataset, the model underestimated harvest rates for 2024, suggesting that the compression effect was underrepresented. Although similar effort compression in response to shortened seasons has been documented for Gulf red snapper (e.g., powers2016estimating, ; powers2018compression, ; topping2019comparison, ; farmer2020forecasting, ), this study provides the first evidence of such effects in another important Gulf fishery. Given these findings, we recommend assessing harvest data for other vulnerable and highly targeted species with recently shortened seasons for evidence of effort compression, to better manage harvest quotas and minimize overfishing risks.
@@ 5 Competing interests
No competing interest is declared.

### v2

@@ 1 Introduction
In this study, we apply a regression framework to forecast harvest for the recreational sector component of the Gulf gag (Mycteroperca microlepis) stock. Gulf gag are highly targeted by recreational anglers (Hyman2024Effort, ) and the gag fishery is currently in a rebuilding plan (Amendment56, ). Recreational catch-per-unit-effort (CPUE) has declined in recent years (GAGSEDAR72, ) and stock status is currently considered impaired (i.e., overfishing is occurring and the stock is overfished; GAGSEDAR72, ; Gag_INTERIM, ; Amendment56, ). In response, the Gulf of Mexico Fishery Management Council (GMFMC) reduced the recreational gag annual catch limit (ACL) by 80% and reduced the season duration from 214 days in 2022 to 49 days in 2023 (Amendment56, ) and 15 days in 2024 (NOAA2024GagSeason, ). Projections of season duration based on harvest rates from previous years underestimated harvest in the substantially shorter season, and the ACL was exceeded in both 2023 (Amendment56, ; NOAA2024GagSeason, ) and with preliminary 2024 estimates, suggesting intensified fishing effort within the shortened time frame. Consequently, more precise projection methodologies that (1) incorporate changes in harvest rates driven by temporal restrictions (i.e., effort compression) and environmental conditions and (2) quantify the risk of exceeding the harvest quota could enhance future fisheries management decisions. We developed a set of statistical models to generate monthly estimates of harvest rates (pounds per day open) using management, socioeconomic, environmental, and seasonal predictors. We also demonstrate this framework may be used to generate pre-season forecasts and in-season updates of when a hypothetical quota will be met.
@@ 2.1 Data sources
We compiled recreational gag harvest data for the for-hire sector from the National Marine Fisheries Service (NMFS) Marine Recreational Information Program (MRIP) For-Hire Telephone Survey (FHTS) combined with data from the Angler Point Access Intercept Survey (APAIS) and private recreational data from the Florida Fish and Wildlife (FWC) State Reef Fish Survey (SRFS). This approach aligns with the data sources currently used for gag management, as both the GMFMC and NOAA NMFS monitor recreational gag landings using SRFS and FHTS/APAIS data.
@@
The State Reef Fish Survey (SRFS) was established in 2015 (originally named the Gulf Reef Fish Survey [GRFS]) to improve harvest estimates for offshore species that are less frequently targeted and caught. The survey focuses on reef-associated species, including gag, red grouper (Epinephelus morio), gray triggerfish (Balistes capriscus), red snapper, vermilion snapper, greater amberjack (Seriola dumerili), hogfish (Lachnolaimus maximus), mutton snapper (Lutjanus analis), and yellowtail snapper (Ocyurus chrysurus). The SRFS estimates CPUE using a combination of data collected on these reef-associated species from the APAIS and data collected from SRFS intercepts, which are conducted at public boat ramps where offshore fishermen are more commonly encountered. The SRFS mail survey is sent to a stratified, random sample of anglers with the free state reef fish angler designation on their Florida saltwater license, asking them to report their fishing activity over the past month. The combination of increased dockside sampling at offshore-access points, a larger sample size, and a more targeted mail survey frame allows the SRFS to produce more precise harvest estimates for reef fish compared to the MRIP survey, which has a broader scope of all saltwater fish species.
@@
Both SRFS and FHS use dockside sampling to collect catch-per-trip data. In these interviews, trained samplers record details from anglers about their completed fishing trips, including the number and species of fish caught, the size of landed fish, and fishing effort (e.g., hours fished). These catch data are then combined with effort estimates from the mail or telephone surveys to calculate total harvest. Statistical methods are applied to account for sampling variability and potential reporting biases, ensuring robust estimates for fisheries management. To standardize harvest rates from SRFS and FHTS in Florida, we aggregated monthly gag harvest estimates across months with varying open seasons. We then divided each monthly estimate by the number of days open to fishing to calculate daily harvest rates. To avoid division by zero in months when harvest was closed, we added one additional day to the denominator.
@@ 2.2 Predictors considered
We considered multiple regulatory, socioeconomic, and environmental predictors to model patterns in recreational gag harvest rates (Table 1). Fundamentally, we expected gag harvest rates to be positively related to whether a month was open to gag harvest, as anglers would presumably target gag when the season is open (Hyman2025Catch, ). However, gag temporal regulations along the Gulf Coast of Florida are complex and have historically differed spatially, requiring careful consideration of how to translate these regulations into relevant predictors. From 2012 to 2022, a four-county special season for gag in Florida applied to Franklin, Wakulla, Jefferson, and Taylor counties. During this period, state waters in these counties were opened for recreational gag harvest from April 1 through June 1 or June 30, depending on the year (Hyman2024Effort, ). This special season allowed for local adjustments to manage the species effectively and support angling opportunities. For all Gulf state and federal waters, gag harvest was generally open from June 1 to December 31. To account for spatial differences in gag temporal restrictions, we created three dummy variables denoting whether the four-county season (denoted “$\text{Special}_{Gag}$”) or the remainder of Gulf state and federal waters (denoted “$\text{Full}_{Gag}$”) was open, or neither (denoted “$\text{Closed}_{Gag}$”). Furthermore, for the hurdle component of our regression model (see Analysis section below), we introduced a fourth binary variable, “$\text{Open}_{Gag}$”, which indicated whether either season was open. This was based on the expectation that the probability of observing zero gag harvested per day would depend on whether gag harvest anywhere was permitted during that period.
@@ 2.3 Analysis
where $y_{t}$ denotes the weight of gag harvested per day open in the $t^{th}$ month-year. Meanwhile, $\mu_{t}$ and $\sigma^{2}_{t}$ denote the mean and variance of the Gamma distribution conditioned on $y_{t}>0$, respectively, and $\theta_{t}$ denotes the binomial probability of observing 0 gag caught in month-year $t$. We use $x$, $z$, and $k$ to refer to predictors included in modeling $\mu_{t}$, $\sigma^{2}_{t}$, and $\theta_{t}$, respectively. Similarly, coefficients corresponding to $\mu_{t}$, $\sigma^{2}_{t}$, and $\theta_{t}$ are denoted $\beta$, $\rho$, and $\lambda$, respectively. Both $\mu_{t}$ and $\sigma^{2}_{t}$ are related to their predictors and associated coefficients through a log-link function, whereas $\theta_{t}$ is related to its predictors and associated coefficients through a logit-link function. All regression coefficients were assigned relatively uninformative, normally-distributed priors with mean 0 and variance 16.
We used the Stan probabilistic programming language for Bayesian modeling (gelman2015stan, ), accessed via R (Rcore2024, ), to fit all models. Each model involved four parallel Markov chains with 5,000 iterations for warm-up and another 5,000 for posterior sampling, yielding 20,000 total draws for inference. Covariates were deemed significant if their 80% confidence intervals excluded 0 (ipccAR6, ; kruschke2021bayesian, ; van2021bayesian, ). All graphics were produced using the ggplot2 package in R (ggplot2, ).
@@ 2.3.2 Cross-validation
We employed extensive cross-validation (CV) to assess the predictive power of our best-performing statistical model. We considered two CV exercises – block-CV and 1-step-ahead-performance-CV (1-SAP-CV) – to separately assess the suitability of our model for pre-season and in-season projections, respectively. For our block-CV exercise, we refit the model to truncated datasets eight times – each with one year withheld – and subsequently compared predictions derived from our models to withheld, observed response variables for all months in the withheld year. As managers would only have information from previous years when making pre-season projections, this cross-validation method was a reasonable approach to assessing model pre-season posterior prediction accuracy. In contrast, nowcasting entails making in-season projections for the current time-step based on previous information both from the current year and previous years (carter2015nowcasting, ). As a result, for 1-SAP CV, for each year, we subsequently refit the model 11 times (once for each month). For each refit, we included months 1 to $m$ in our training set and then compared posterior predictive distributions to observed data in month $m+1$.
@@ 2.4 Comparison to current methods
After establishing the posterior predictive accuracy of our models, we conducted simulations to predict the 2025 gag season duration using 1,000 posterior scans from our best-fitting model. To begin, we calculated the adjusted annual catch target (ACT) for 2025 by accounting for the 2024 overage penalty—defined as the amount of gag harvested in 2024 (249,000 lbs) that exceeded the 2024 ACL (163,000 lbs). This overage was subtracted from the 2025 ACT (319,000 lbs) specified in the rebuilding plan (Amendment56, ; NOAA2024GagSeason, ; ECFR2025, ):
@@
To develop our simulations, we constructed a design matrix where all predictors, except for gag temporal regulations and seasonal variables, were fixed at their 2024 values. We began by setting the season duration to $d=1$ day, with a starting date of September 1st. Months without any days open to harvest were assigned the “Closed” temporal regulation treatment, whereas months with days open were assigned the “Full” treatment. Using our best-fitting model, we simulated posterior-predictive distributions of monthly harvest rates for 2025 across 1,000 posterior scans ($s$). For each scan, monthly harvest rates were multiplied by the number of days open to harvest within the corresponding month to derive posterior predictions of monthly harvest. To estimate total harvest for each scan ($s$), we summed harvest estimates across all months. The process was repeated iteratively, incrementing the season duration ($d$) by one day at a time, until $d=121$, which corresponded to the maximum possible season duration ending December 31st. For each season duration ($d=1,\dots,121$), we calculated the median and 80% posterior predictive interval of total harvest. Additionally, for each season duration, we estimated the probability that the ACT would be met or exceeded as the proportion of posterior predictive scans whose total harvest met or exceeded the ACT. Finally, we compared these projections to those using the historic average harvest rate for September.
@@ 3.1 Model selection and diagnostics
Time-series plots comparing observed and predicted values, along with posterior predictive checks, indicated that model $g_{2}$ provided a good fit to the observed data. The posterior predictive median and 80% CI of monthly estimates aligned closely with in-sample estimates of monthly gag harvest rates (Fig. 1a). Additionally, diagnostic scatter-plots of median predicted versus observed gag harvest rates and comparisons of empirical distributions to posterior predictive distributions derived from model $g_{2}$ demonstrated satisfactory model performance (Figs. 1b and 1c).
@@ 3.2 Predictors of monthly gag harvest rates
Gag temporal regulations, the annual trend, and seasonal harmonic terms all significantly explained variation in mean (nonzero) recreational gag harvest rates (Table 3.2). Conditioned on a positive outcome and relative to a closed season, months when gag was open to the entire Florida Gulf Coast (FullGag) resulted in higher harvest rates, while months open only for the four-county season (SpecialGag) resulted in lower harvest rates. The duration of the gag season had a clear negative effect on harvest rates, such that the harvest rate was lower during longer gag seasons (Fig. 2). The annual trend was positive and significant, indicating that after accounting for other predictors, gag harvest rates appeared to be increasing over time. Notably, after accounting for the effects of other predictors, past harvest did not significantly explain residual variation in gag harvest rates.
@@
The hurdle parameter, $\theta$ was strongly and negatively impacted by whether gag was open to harvest (either in the four-county region or across the Gulf Coast of Florida). When the season was completely closed to gag harvest, the probability of observing 0 gag harvested in a given month was 0.90 (80% CI: 0.84 – 0.95). However, this probability decreased to nearly 0 when gag was open to harvest. Meanwhile, both sinusoidal harmonic terms significantly explained $\sigma^{2}$, suggesting that model uncertainty varied seasonally.
@@ 4.1 Predictors of gag harvest
A key finding of this study is the pronounced effort compression observed in the recreational Gulf gag fishery. While previous research did not identify a relationship between gag season duration and recreational reef fishing effort (e.g., Hyman2024Effort, ), our model revealed a strongly negative effect on gag harvest rates, which increased non-linearly as season duration decreased. As noted by Hyman2024Effort , a lack of contrast in season duration previously may have limited the detection of this effect. The recreational gag fishing seasons have only recently been significantly shortened, with the 2023 and 2024 seasons being the shortest on record. This is supported by our CV results: when 2024 data were included in model training, the model accurately predicted effort compression and correctly estimated withheld gag harvest rates for 2023 (Fig. 4). However, when 2024 data were excluded from the training dataset, the model underestimated harvest rates for 2024, suggesting that the compression effect was underrepresented. Although similar effort compression in response to shortened seasons has been documented for Gulf red snapper (e.g., powers2016estimating, ; powers2018compression, ; topping2019comparison, ; farmer2020forecasting, ), this study provides the first evidence of such effects in another important Gulf fishery. Given these findings, we recommend assessing harvest data for other vulnerable and highly targeted species with recently shortened seasons for evidence of effort compression, to better manage harvest quotas and minimize overfishing risks.
@@ 5 Competing interests
No competing interests are declared.